### v1

# label*label*footnotetext: Corresponding authors

Abstract
This note aims to present novel positive linear operators involving the Wright function. Furthermore, the present research established the moments of these newly defined operators and estimated the convergence rate using the classical modulus of continuity. Additionally, the convergence rate in the Lipschitz spaces, their $A$-statistical convergence property, and $(\lambda,\gamma)-$ statistical convergence properties have been covered.
keywords:
Wright function; hypergeometric function; Bessel function; $A$-statistical convergence
2000 Mathematics Subject Classification: primary 41A25, 41A30, 41A36.
journal: arXiv
1 Introduction
A new generalization of Bernstein operators was introduced periodically in approximation theory (see totik1994approximation ; lewanowicz2004generalized ; sinha2009inverse ; gupta2013approximation ; gupta2021operators ; mishra2014generalized ). Many positive linear operators are derived from special functions. For instance, the operators were introduced and studied in ozarslan2013astatistical using the Mittag-Leffler function, the operators were studied in gupta2020convergence with the help of gamma function, and using Laguerre polynomials, the positive linear operators introduced by Gupta gupta2024new . This note aims to merge two mathematical topics-approximation theory and special functions.
The special function
$\phi(\rho,\beta;z)=\phi_{\rho,\beta}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{k!\Gamma(\rho k+\beta)},~{}~{}~{}(\rho>-1,\beta,z\in\mathbb{C})$ (1.1)
named after the British mathematician E. M. Wright, has appeared for the first time in the case $\rho>0$ in connection with his investigations in the asymptotic theory of partitions in wright1933coefficients .
Note that $\phi(1,1;z)=I_{0}(2\sqrt{z})$, where $I_{n}(z)=i^{-n}J_{n}(iz)=\dfrac{(z/2)^{n}}{\Gamma(1+n)}{}_{0}F_{1}\left(-;1+n;\dfrac{z^{2}}{4}\right)$, $n$ not a negative integer. The function $I_{n}(z)$ is called the modified Bessel function of the first kind of index $n$. Here, $J_{n}$ is Bessel function and ${}_{0}F_{1}$ is hypergeometric function. The modified function $I_{n}$ is related to $J_{n}$ in much the same way that the hyperbolic function is related to the trigonometric function.
Also, note that
$\phi_{1,2}(z)=\dfrac{1}{\sqrt{z}}I_{1}\left(2\sqrt{z}\right);~{}~{}~{}~{}~{}\phi_{1,m}(z)=\left(\dfrac{1}{\sqrt{z}}\right)^{m/2}I_{m-1}\left(2\sqrt{z}\right).$
It was demonstrated in gray1895treatise ; bateman1953higher that, the modified Bessel function $I_{n}(z)$ of the first kind does not have zeros in the right half-plane. khanmamedov2010zeros contains an analysis of $n$-zeros of the function $I_{n}(z)$ in the left half-plane. Since $I_{n}(z)$ has no zeros in $[0,\infty)$, the Wright function $\phi_{1,m}(z)$ also has no zeros in $[0,\infty)$.
Let $\beta>1$ be fixed. We provide a novel class of positive linear operators involving the Wright function for every $n\in\mathbb{N}$ as
$W_{n}^{(\beta)}(f;x)=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}f\left(\dfrac{k}{n}\right)\dfrac{(nx)^{k}}{k!\Gamma(k+\beta)},$ (1.2)
where $f\in E:=\left\{f\in C\left[0,\infty\right):\lim_{x\to\infty}\dfrac{f(x)}{1+x^{2}}\textrm{ is finite }\right\}$ and $C\left[0,\infty\right)$ denote the space of continuous functions defined on $[0,\infty)$. Recall that the Banach lattice $E$ is endowed with the norm
$\|f\|_{2}:=\sup_{x\in[0,\infty)}\dfrac{|f(x)|}{1+x^{2}}.$
One can note that the operators $W_{n}^{(\beta)}$ defined in (1.2) are linear and positive. We called it the Wright operators.
Research on approximating continuous signals using a sequence of positive linear operators is ongoing. We provide other findings along the same lines in this note, but with the aid of recently established positive linear operators $W_{n}^{(\beta)}$. There is no prior literature that introduced these operators (1.2). We estimate the moments and central moments of the operators (1.2) up to the fourth order in the second section. The rate of convergence of these positive linear operators is discussed in the next section. In the same section 3, we have proved that $W_{n}^{(\beta)}$ maps $E$ into itself. We have also determined the rate of convergence for these operators by utilizing the modulus of continuity. For the operators $W_{n}^{(\beta)}$, we have established a statistical Voronovskaya-type theorem in section 4. A weighted $\lambda\beta$-statistical approximation theorem for the Wright operators (1.2) has been established in section 5. It should be noted that depending on the choice of $\lambda$ and $\gamma$, $\lambda\gamma$-statistical convergence is a non-trivial extension of ordinary and statistical convergence. Some extensions of the Wright operators are shown in the last part to estimate the functions in $L_{p}$ space for the next research.
2 Some Lemmas
To analyze approximation features of the operators (1.2), a few inequalities for the Wright function are required.
Lemma 1
(mehrez2017functional, , Theorem 6.1) Let $\alpha,\beta>0$. Then the following assertions are true:
$\Gamma(\beta+\alpha)\phi_{\alpha,\beta+\alpha}(z)\leq\Gamma(\beta)\phi_{\alpha,\beta}(z),\textrm{ for any }z>0.$ (2.1)
Direct calculations allow one to declare the following lemma:
Lemma 2
Let $\phi_{x}^{2}\left(t\right)=\left(t-x\right)^{2}$, for each $x\geq 0$, $\beta>1$ and $n\in\mathbb{N}$, we have
  1. 1.
    $W_{n}^{(\beta)}(1;x)=1;$
  2. 2.
    $\left|W_{n}^{(\beta)}(t;x)-x\right|\leq 0;$
  3. 3.
    $\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|\leq\dfrac{x}{n\beta};$
  4. 4.
    $\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|\leq\dfrac{3x^{2}}{n\beta(\beta+1)}+\dfrac{x}{n^{2}\beta};$
  5. 5.
    $\left|W_{n}^{(\beta)}(t^{4};x)-x^{4}\right|\leq\dfrac{6x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{7x^{2}}{n^{2}\beta(\beta+1)}+\dfrac{x}{n^{3}\beta}.$
Proof: Since,
$\sum_{k=0}^{\infty}\frac{(nx)^{k}}{k!\Gamma(k+\beta)}=\phi_{1,\beta}(nx),$
one can have $W_{n}^{(\beta)}(1,x)=1$. Using Lemma 1, $\Gamma(\beta+1)\phi_{1,\beta+1}\left(nx\right)\leq\Gamma(\beta)\phi_{1,\beta}\left(nx\right)$ for any $x\in[0,\infty)$ and $n\in\mathbb{N}$, we get
$W_{n}^{(\beta)}(t;x)$ $=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=1}^{\infty}\dfrac{1}{n}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}$
$=\dfrac{x}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{(nx)^{k}}{k!\Gamma(k+\beta+1)}=x\dfrac{\phi_{1,\beta+1}(nx)}{\phi_{1,\beta}(nx)}\leq\dfrac{x}{\beta}\leq x\textrm{ for }\beta>1.$
Again applying Lemma 1, $\Gamma(\beta+2)\phi_{1,\beta+2}\left(nx\right)\leq\Gamma(\beta+1)\phi_{1,\beta+1}\left(nx\right)\leq\Gamma(\beta)\phi_{1,\beta}\left(nx\right)$ for any $x\in[0,\infty)$ and $n\in\mathbb{N}$, we get
$W_{n}^{(\beta)}(t^{2};x)$ $=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=1}^{\infty}\dfrac{(k-1)+1}{n^{2}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}$
$=\dfrac{x^{2}}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{(nx)^{k}}{k!\Gamma(k+\beta+2)}+\dfrac{x}{n\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{(nx)^{k}}{k!\Gamma(k+\beta+1)}$
$=x^{2}\dfrac{\phi_{1,\beta+2}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}+\dfrac{x}{n}\dfrac{\phi_{1,\beta+1}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}\leq\left|\dfrac{x^{2}\Gamma(\beta)}{\Gamma(\beta+2)}\right|+\left|\dfrac{x\Gamma(\beta)}{n\Gamma(\beta+1)}\right|=\dfrac{x^{2}}{\beta(\beta+1)}+\dfrac{x}{\beta n}.$
Since $\beta>1$,
$\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|\leq\dfrac{x}{\beta n}.$
Now,
$W_{n}^{(\beta)}(t^{3};x)$ $=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=1}^{\infty}\dfrac{k^{2}}{n^{3}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}$
$=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=2}^{\infty}\dfrac{k+1}{n^{3}}\dfrac{(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{1}{n^{3}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=3}^{\infty}\dfrac{1}{n^{3}}\dfrac{(nx)^{k}}{(k-3)!\Gamma(k+\beta)}+\sum_{k=2}^{\infty}\dfrac{3}{n^{3}}\dfrac{(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{1}{n^{3}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $x^{3}\dfrac{\phi_{1,\beta+3}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}+3x^{2}\dfrac{\phi_{1,\beta+2}\left(nx\right)}{n\phi_{1,\beta}\left(nx\right)}+x\dfrac{\phi_{1,\beta+1}\left(nx\right)}{n^{2}\phi_{1,\beta}\left(nx\right)}$
$\leq$ $\left|\dfrac{x^{3}\Gamma(\beta)}{\Gamma(\beta+3)}\right|+\left|\dfrac{3x^{2}\Gamma(\beta)}{n\Gamma(\beta+2)}\right|+\left|\dfrac{x\Gamma(\beta)}{n^{2}\Gamma(\beta+1)}\right|=\dfrac{x^{3}}{\beta(\beta+1)(\beta+2)}+\dfrac{3x^{2}}{n\beta(\beta+1)}+\dfrac{x}{\beta n^{2}}.$
Since, $\beta>1$,
$\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|\leq\dfrac{3x^{2}}{n\beta(\beta+1)}+\dfrac{x}{\beta n^{2}}.$
Similarly,
$W_{n}^{(\beta)}(t^{4};x)$ $=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=1}^{\infty}\dfrac{k^{3}-1+1}{n^{4}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}$
$=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=2}^{\infty}\dfrac{k^{2}+k+1}{n^{4}}\dfrac{(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{1}{n^{4}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=2}^{\infty}\dfrac{k^{2}-4+4}{n^{4}}\dfrac{(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=2}^{\infty}\dfrac{k+1}{n^{4}}\dfrac{(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{1}{n^{4}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $\dfrac{1}{n^{4}\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=3}^{\infty}\dfrac{(k+2)(nx)^{k}}{(k-3)!\Gamma(k+\beta)}+\sum_{k=2}^{\infty}\dfrac{4(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=2}^{\infty}\dfrac{(k-2+3)(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $\dfrac{1}{n^{4}\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=4}^{\infty}\dfrac{(nx)^{k}}{(k-4)!\Gamma(k+\beta)}+\sum_{k=2}^{\infty}\dfrac{7(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=3}^{\infty}\dfrac{6(nx)^{k}}{(k-3)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $x^{4}\dfrac{\phi_{1,\beta+4}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}+6x^{3}\dfrac{\phi_{1,\beta+3}\left(nx\right)}{n\phi_{1,\beta}\left(nx\right)}+7x^{2}\dfrac{\phi_{1,\beta+2}\left(nx\right)}{n^{2}\phi_{1,\beta}\left(nx\right)}+x\dfrac{\phi_{1,\beta+1}\left(nx\right)}{n^{3}\phi_{1,\beta}\left(nx\right)}$
$\leq$ $\left|\dfrac{x^{4}\Gamma(\beta)}{\Gamma(\beta+4)}\right|+\left|\dfrac{6x^{3}\Gamma(\beta)}{n\Gamma(\beta+3)}\right|+\left|\dfrac{7x^{2}\Gamma(\beta)}{n^{2}\Gamma(\beta+2)}\right|+\left|\dfrac{x\Gamma(\beta)}{n^{3}\Gamma(\beta+1)}\right|$
$=$ $\dfrac{x^{4}}{\beta(\beta+1)(\beta+2)(\beta+3)}+\dfrac{6x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{7x^{2}}{n^{2}\beta(\beta+1)}+\dfrac{x}{n^{3}\beta}.$
Since, $\beta>1$,
$\left|W_{n}^{(\beta)}(t^{4};x)-x^{4}\right|\leq\dfrac{6x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{7x^{2}}{n^{2}\beta(\beta+1)}+\dfrac{x}{n^{3}\beta}.$
Lemma 3
Let $\phi_{x}^{i}\left(t\right)=\left(t-x\right)^{i}$, $i\in\mathbb{N}$, for each $x\geq 0$, $\beta>1$ and $n\in\mathbb{N}$, we have
  1. 1.
    $W_{n}^{(\beta)}(\phi_{x}^{1};x)=0$
  2. 2.
    $W_{n}^{(\beta)}(\phi_{x}^{2};x)\leq\dfrac{x}{n\beta}$
  3. 3.
    $W_{n}^{(\beta)}(\phi_{x}^{3};x)\leq\dfrac{3x^{2}(\beta+2)}{n\beta(\beta+1)}+\dfrac{x}{n^{2}\beta}.$
  4. 4.
    $W_{n}^{(\beta)}(\phi_{x}^{4};x)\leq\dfrac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{n\beta(\beta+1)(\beta+2)}+\dfrac{4x^{2}\beta+11}{n^{2}\beta(\beta+1)}+\dfrac{x}{n^{3}\beta}.$
Proof: Using Lemma 2 and simple computation gives $W_{n}^{(\beta)}(\phi_{x}^{1};x)=0$ and
$W_{n}^{(\beta)}(\phi_{x}^{2};x)\leq\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|+2x\left|W_{n}^{(\beta)}(t;x)-x\right|+x^{2}\left|W_{n}^{(\beta)}(1;x)-1\right|\leq\dfrac{x}{n\beta}.$
Further,
$W_{n}^{(\beta)}(\phi_{x}^{3};x)$ $\leq\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|+3x\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|+3x^{2}\left|W_{n}^{(\beta)}(t;x)-x\right|+x^{3}\left|W_{n}^{(\beta)}(1;x)-1\right|$
$\leq\dfrac{3x^{2}}{n\beta(\beta+1)}+\dfrac{x}{n^{2}\beta}+3x\left(\dfrac{x}{n\beta}\right)=\dfrac{3x^{2}(\beta+2)}{n\beta(\beta+1)}+\dfrac{x}{n^{2}\beta}.$
Finally, we have
$W_{n}^{(\beta)}(\phi_{x}^{4};x)$ $\leq\left|W_{n}^{(\beta)}(t^{4};x)-x^{4}\right|+4x\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|$
$+6x^{2}\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|+4x^{3}\left|W_{n}^{(\beta)}(t;x)-x\right|+x^{4}\left|W_{n}^{(\beta)}(1;x)-1\right|$
$\leq\dfrac{6x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{7x^{2}}{n^{2}\beta(\beta+1)}+\dfrac{x}{n^{3}\beta}+4x\left(\dfrac{3x^{2}}{n\beta(\beta+1)}+\dfrac{x}{n^{2}\beta}\right)+6x^{2}\left(\dfrac{x}{n\beta}\right)$
$=\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{n\beta(\beta+1)(\beta+2)}+\frac{4x^{2}\beta+11}{n^{2}\beta(\beta+1)}+\frac{x}{n^{3}\beta}.$
3 Rate of Convergence
In this section, the rate of convergence of $W_{n}^{(\beta)}$ is discussed. The following lemma proves that $W_{n}^{(\beta)}$ maps $E$ into itself.
Lemma 4
Let $\beta>1$ be fixed, then there exists a constant $M(\beta)$ such that,
$\omega(x)W_{n}^{(\beta)}\left(\dfrac{1}{\omega},x\right)\leq M(\beta)$
holds for all $x\in[0,\infty)$, $n\in\mathbb{N}$ and $\omega(x)=\dfrac{1}{1+x^{2}}$. Furthermore, for all $f\in E$, we have
$\|D_{n}^{(\beta)}(f,\cdot)\|_{2}\leq M(\beta)\|f\|_{2}$
Proof: From Lemma 2, we have
$\omega(x)W_{n}^{(\beta)}\left(\dfrac{1}{\omega},x\right)$ $=$ $\dfrac{1}{1+x^{2}}\left[W_{n}^{(\beta)}\left(1,x\right)+W_{n}^{(\beta)}\left(t^{2},x\right)\right]$
$\leq$ $\dfrac{1}{1+x^{2}}\left[1+x^{2}+\dfrac{x^{2}}{\beta(\beta+1)}+\dfrac{x}{\beta n}\right]$
$\leq$ $M(\beta)$
This follows, following inequality
$\omega(x)\left|W_{n}^{(\beta)}\left(f,x\right)\right|=\omega(x)\left|W_{n}^{(\beta)}\left(\omega\dfrac{f}{\omega},x\right)\right|\leq\|f\|_{2}\omega(x)W_{n}^{(\beta)}\left(\dfrac{1}{\omega},x\right)\leq M(\beta)\|f\|_{2}$
Taking supremum over $x\in[0,\infty)$ in the above inequality, gives the result. $\blacksquare$
Now, recall that the usual modulus of continuity of $f$ on the closed interval $[0,B]$ is defined by
$\omega_{B}(f,\delta)=\sup\{|f(t)-f(x)|:|t-x|\leq\delta,~{}~{}x,t\in[0,B]\}$
It is well known that, for a function $f\in E$, we have $\lim_{\delta\to\infty}\omega_{B}(f,\delta)=0$. The next theorem gives the rate of convergence of the operators $W_{n}^{(\beta)}(f,x)$, for all $f\in E$.
Theorem 1
Let $\beta>1$, $f\in E$ and $\omega_{B+1}(f,\delta)$, $(B>0)$ be its modulus of continuity on the finite interval $[0,B+1]\subset[0,\infty)$, then
$\|W_{n}^{(\beta)}\left(f,\cdot\right)-f(\cdot)\|_{C[0,B]}\leq M_{f}(\beta,B)\delta_{n}(\beta,B)+2\omega_{B+1}\left(f,\delta_{n}^{1/2}(\beta,B)\right)$
where $\delta_{n}(\beta,B)=\dfrac{B}{n\beta}$ and $M_{f}(\beta,B)$ is an absolute constant depending on $f$, $\beta$ and $B$.
Proof: Let $\beta>0$ be fixed. For $x\in[0,B]$ and $t\leq B+1$, we have following well-known inequality
$|f(t)-f(x)|\leq\omega_{B+1}(f,|t-x|)\leq\left(1+\dfrac{|t-x|}{\delta}\right)\omega_{B+1}(f,\delta)$ (3.1)
where $\delta>0$. Now, for $x\in[0,B]$ and $t>B+1$, using the fact that $t-x>1$, we have
$|f(t)-f(x)|$$\leq$ $A_{f}(1+x^{2}+t^{2})$ (3.2)
$\leq$ $A_{f}(2+3x^{2}+2(t-x)^{2})$
$\leq$ $6A_{f}(1+B^{2})(t-x)^{2}$
Using (3.1) and (3.2), we get for all $x\in[0,B]$ and $t\geq 0$, we get
$|f(t)-f(x)|$$\leq$ $6A_{f}(1+B^{2})(t-x)^{2}+\left(1+\dfrac{|t-x|}{\delta}\right)\omega_{B+1}(f,\delta)$ (3.3)
Therefore,
$\left|W_{n}^{(\beta)}\left(f,x\right)-f(x)\right|\leq 6A_{f}(1+B^{2})W_{n}^{(\beta)}\left(\phi_{x}^{2},x\right)+\left(1+\dfrac{W_{n}^{(\beta)}\left(|t-x|,x\right)}{\delta}\right)\omega_{B+1}(f,\delta)$
Applying Cauchy-Schwarz inequality and Lemma 2, we get
$\left|W_{n}^{(\beta)}\left(f,x\right)-f(x)\right|$$\leq$ $6A_{f}(1+B^{2})W_{n}^{(\beta)}\left(\phi_{x}^{2},x\right)+\left(1+\dfrac{\left[W_{n}^{(\beta)}\left(\phi_{x}^{2},x\right)\right]^{1/2}}{\delta}\right)\omega_{B+1}(f,\delta)$
$\leq$ $6A_{f}(1+B^{2})\left(\dfrac{B}{n\beta}\right)+\left(1+\dfrac{\left[\dfrac{B}{n\beta}\right]^{1/2}}{\delta}\right)\omega_{B+1}(f,\delta)$
$\leq$ $M_{f}(\beta,B)\delta_{n}^{2}(\beta,B)+2\omega_{B+1}\left(f,\left(\delta_{n}(\beta,B)\right)^{1/2}\right)$
where $M_{f}(\beta,B)=6A_{f}(1+B^{2})$ and $\delta_{n}(\beta,B)=\dfrac{B}{n\beta}$. Hence the proof. $\blacksquare$
4 $A$-statistical Convergence
In this section, first we discussed some definitions and notations on the concept of $A$-statistical convergence. Let $A=\left(a_{nk}\right)$, $\left(n,k\in\mathbb{N}\right)$, be a non-negative, infinite summability matrix. For a given sequence $x:=(x_{k})$, the $A$-transform of $x$ denoted by $Ax:\left((Ax)_{n}\right)$ is defined as
$\left(Ax\right)_{n}=\sum_{k=1}^{\infty}a_{nk}x_{k},$
provided the series converges for each $n$. $A$ is said to be regular if $\lim_{n}(Ax)_{n}=L$ whenever $\lim_{n}x_{n}=L$. We say that, the sequence $x=(x_{n})$ is $A$-statistically convergent to $L$ and write $st_{A}-\lim_{n}x_{n}=L$ if for every $\epsilon>0$, $\lim_{n}\sum_{k:|x_{k}-L|\geq\epsilon}a_{nk}=0$.
Replacing $A$ by $C_{1}$, the Cesáro matrix of order one, the $A$-statistical convergence reduces to statistical convergence. Similarly, if we take $A=I$, the identity matrix, then $A$-statistical convergence coincides with ordinary convergence. The statistical convergence of various types of operators has been studied by several researchers (see mishra2017statistical ; duman2008statistical ; gadjiev2002some ; duman2006statistical ; gupta2009statistical ; ispir2008statistical ; gupta2014convergence ; gupta2021modifications ; gupta2021higher ; mursaleen2009invariant ; gurhan2015durrmeyer ; garrancho2019general ; chandra2022approximation ; agrawal2022stancu ; maurya2018approximation ). Now, we prove weighted Korovkin theorem via $A$-statistical convergence.
Here, we recall the weighted Korovkin type approximation theorem for the A-statistical convergence was given by Duman and Orhan in duman2004statistical .
Theorem 2
duman2004statistical Let $A$ be a non-negative regular summability matrix and let $\bar{\rho}_{1}$; $\bar{\rho}_{2}$ weight functions such that
$\lim_{|x|\to\infty}\dfrac{\bar{\rho}_{1}(x)}{\bar{\rho}_{2}(x)}=0.$
Assume that $(T_{n})_{n\geq 1}$ is a sequence of positive linear operators from $C_{\bar{\rho}_{1}}(\mathbb{R})$ into $B_{\bar{\rho}_{2}}(\mathbb{R})$, One has
$st_{A}-\lim_{n}\|T_{n}f-f\|_{\bar{\rho}_{2}}=0,$
for all $f\in C_{\bar{\rho}_{1}}(\mathbb{R})$ if and only if
$st_{A}-\lim_{n}\|T_{n}F_{v}-F_{v}\|_{\bar{\rho}_{2}}=0,\textrm{ for all }v=0,1,2,$
where
$F_{v}(x)=\dfrac{x^{v}\bar{\rho}_{1}(x)}{1+x^{2}},v=0,1,2.$
By using this theorem the following Korovkin type theorem can be proved for $\left(W_{n}^{(\beta)}\right)$.
Theorem 3
Let $A=(a_{nk})$ be a non-negative regular summability matrix, $\beta>1$ be fixed and $x\in[0,\infty)$, then for all $f\in E$, we have
$st_{A}-\lim_{n}\|W_{n}^{(\beta)}\left(f,\cdot\right)-f\|_{2}=0$
Proof: From (duman2004statistical, , p. 191, Th. 3), it is sufficient to show that $st_{A}-\lim_{n}\|W_{n}^{(\beta)}\left(t^{i},\cdot\right)-x^{i}\|_{2}=0$, where $i=0,1,2$.
In view of Lemma 2, it follows that
$st_{A}-\lim_{n}\|W_{n}^{(\beta)}\left(1,\cdot\right)-1\|_{2}=0$
and
$st_{A}-\lim_{n}\|W_{n}^{(\beta)}\left(t,\cdot\right)-x\|_{2}=0.$
Now,
$\|W_{n}^{(\beta)}\left(t^{2},\cdot\right)-x^{2}\|_{2}\leq\sup_{x\geq 0}\left(\dfrac{x}{n\beta(1+x^{2})}\right)\leq\dfrac{1}{n\beta}$
Given $r>0$, choose $\epsilon>0$ such that $\epsilon<r$. For fixed $\beta>1$, define the following sets:
$U:=\left\{n:\|W_{n}^{(\beta)}\left(t,\cdot\right)-x\|_{2}\geq\epsilon\right\}$
$U_{1}:=\left\{n:\frac{1}{n\beta}\geq\dfrac{\epsilon}{2}\right\}$
Then it is clear that $U\subset U_{1}$, this gives
$\sum_{k\in U}a_{jk}\leq\sum_{k\in U_{1}}a_{jk}$ (4.1)
Letting $j\to\infty$ in (4.1), we have $\lim_{j}\sum_{k\in U}a_{jk}=0$. This proves that $st_{A}-\lim_{n}\dfrac{1}{n\beta}=0$, this also implies
$st_{A}-\lim_{n}\omega_{B+1}\left(f,\sqrt{\dfrac{1}{n\beta}}\right)=0$
Using theorem 1, we get desired result. $\blacksquare$
Lemma 5
Let $A=(a_{in})$ be a non-negative regular summability matrix, then we have
$st_{A}-\lim_{n\to\infty}nW_{n}^{(\beta)}\left((t-x)^{4},x\right)=\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{\beta(\beta+1)(\beta+2)}$
uniformly with respect to $x\in[0,b]$ with $b>0$.
Proof: Note that
$W_{n}^{(\beta)}\left((t-x)^{4},x\right)$$\leq$ $\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{n\beta(\beta+1)(\beta+2)}+\frac{4x^{2}\beta+11}{n^{2}\beta(\beta+1)}+\frac{x}{n^{3}\beta}.$
This gives
$\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{\beta(\beta+1)(\beta+2)}\right|\leq\dfrac{h_{1}(\beta)}{n}x^{2}+\dfrac{h_{2}(\beta)}{n^{2}}x$
where $h_{1}$ and $h_{2}$ are some functions of $\beta$.
For $x\in[0,b]$, we have
$\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{\beta(\beta+1)(\beta+2)}\right|\leq B\left\{\dfrac{1}{n}+\dfrac{1}{n^{2}}\right\}$ (4.2)
where $B=\max\{h_{1}(\beta)b^{2},h_{2}(\beta)b\}$. Now, for a given $\epsilon>0$, define the following sets:
$D$ ::=:= $\left\{n:\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{\beta(\beta+1)(\beta+2)}\right|\geq\epsilon\right\}$
$D_{1}$ ::=:= $\left\{n:\dfrac{1}{n}\geq\dfrac{\epsilon}{2B}\right\}$
$D_{2}$ ::=:= $\left\{n:\dfrac{1}{n}\geq\sqrt{\dfrac{\epsilon}{2B}}\right\}$
Hence, by inequality (4.2), we see that $D\subset D_{1}\cup D_{2}$. Then for any $j\in\mathbb{N}$, we have
$\sum_{n\in D}a_{jn}\leq\sum_{n\in D_{1}}a_{jn}+\sum_{n\in D_{2}}a_{jn}.$ (4.3)
Taking limit as $j\to\infty$ on the both sides of above inequality and using the fact that $st_{A}-\lim_{n\to\infty}\left(\dfrac{1}{n}\right)=0$, we conclude that
$\lim_{j\to\infty}\sum_{n\in D}a_{jn}=0.$
Hence the result. $\blacksquare$
Theorem 4 (Statistical Voronovskaya-type theorem for the operator $W_{n}^{(\beta)}$)
Let $A=(a_{jn})$ be a nonnegative regular summability matrix, $\beta>1$ then for every $f\in E$ with $f^{\prime},f^{\prime\prime}\in E$, we have
$st_{A}-\lim_{n\to\infty}n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)=\dfrac{x}{2\beta}f^{\prime\prime}(x)$
uniformly with respect to $x\in[0,b]$ with $b>0$.
Proof. Let $f,f^{\prime},f^{\prime\prime}\in E$ and $x\in[0,b]$. Define the function $\Phi_{x}$ by
$\Phi_{x}(t)=\left\{\begin{array}[]{ll}\dfrac{f(t)-f(x)-(t-x)f^{\prime}(x)-\frac{1}{2}(t-x)^{2}f^{\prime\prime}(x)}{(t-x)^{2}},&\textrm{ if }t\neq x\\ 0,&\textrm{ if }t=x\end{array}\right.$
Then, it is clear that $\Phi_{x}(x)=0$. Also, observe that the function $\Phi_{x}(\cdot)$ belongs to $E$. Hence, by Taylor’s theorem, we get
$f(t)=f(x)+(t-x)f^{\prime}(x)+\dfrac{(t-x)^{2}}{2}f^{\prime\prime}(x)+(t-x)^{2}\Phi_{x}(t)$
Now the definition of the operators (1.2) implies that
$W_{n}^{(\beta)}(f,x)-f(x)=f^{\prime}(x)W_{n}^{(\beta)}(t-x,x)+\dfrac{1}{2}f^{\prime\prime}(x)W_{n}^{(\beta)}((t-x)^{2},x)+W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)$
Therefore, using lemma 2, we have
$\left|n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)-\dfrac{x}{\beta}f^{\prime\prime}(x)\right|$$\leq$ $n\left|W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)\right|$ (4.4)
Applying the Cauchy-Schwarz inequality to the second term on the right-hand side of (4.4), then we see that
$\left|W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)\right|\leq\sqrt{W_{n}^{(\beta)}((t-x)^{4},x)}\sqrt{W_{n}^{(\beta)}(\Phi_{x}^{2}(t),x)}$
this yields
$n\left|W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)\right|\leq\sqrt{n^{2}W_{n}^{(\beta)}((t-x)^{4},x)}\sqrt{W_{n}^{(\beta)}(\Phi_{x}^{2}(t),x)}$ (4.5)
Let $\eta_{x}(t):=\Phi_{x}^{2}(t)$. In this case, observe that $\eta_{x}(x)=0$ and $\eta_{x}(\cdot)\in E$. Then it follows from Theorem 3 that
$st_{A}-\lim_{n\to\infty}W_{n}^{(\beta)}(\Phi_{x}^{2}(t),x)=st_{A}-\lim_{n\to\infty}W_{n}^{(\beta)}(\eta_{x}(t),x)=\eta_{x}(x)=0$ (4.6)
uniformly with respect to $x\in[0,b]$. Now considering (4.5) and (4.6), and also Lemma 5, we immediately lead to
$st_{A}-\lim_{n\to\infty}n\left(W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)\right)=0$ (4.7)
uniformly with respect to $x\in[0,b]$. Using (4.4) to (4.7) and also considering $st_{A}-\lim_{n\to\infty}\dfrac{1}{n}=0$, we have
$st_{A}-\lim_{n\to\infty}n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)=\dfrac{x}{\beta}f^{\prime\prime}(x)$
uniformly with respect to $x\in[0,b]$. $\blacksquare$
5 $\lambda\gamma$-statistical convergence
Recently, $\lambda\gamma$-statistical convergence of linear positive operators drawn the attention of several researchers (see karaisa2017alphabeta ; bozma2022approximation ; karaisaa2017alphabeta ; gogoi2021weighted ; karakaya2015korovkin ; braha2020korovkin ; abay2018alphabeta ; aktuuglu2016weighted ). Let $K$ be a set of natural numbers, then the density of $K$ is denoted by $\delta(K)$ and defined as
$\delta(K)=\lim_{n\to\infty}\dfrac{|K_{n}|}{n}$ (5.1)
where $K_{n}:=\{k\leq n:k\in K\}$ and $|S|$ represents the cardinality of $S$. Using the density function given in (5.1), Now, we introduced the concept of statistical convergence defined as fast1951convergence . A sequence $x_{n}$ is called statistically convergent to $l$ and denoted by $st-\lim_{n\to\infty}x_{n}=l$ if for every $\epsilon>0$, $\delta(K_{\epsilon})=0$ where $K_{\epsilon}:=\{k\leq n:|x_{k}-l|\geq\epsilon\}$.
Recently, the concept of $\lambda\gamma$-statistical convergence introduced by aktuuglu2014korovkin as follows:
Assume that $\lambda(n)$ and $\gamma(n)$ are two sequences of positive numbers satisfying the following conditions:
  1. $(C_{1})$
    $\lambda$ and $\gamma$ are both non-decreasing,
  2. $(C_{2})$
    $\gamma(n)\geq\lambda(n)$ for all $n\in\mathbb{N}$,
  3. $(C_{3})$
    $\gamma(n)-\lambda(n)\to\infty$ as $n\to\infty$,
and $\Lambda$ denotes the set of pairs $(\lambda,\gamma)$ satisfying $(C_{1})$, $(C_{2})$ and $(C_{3})$.
For each pair $(\lambda,\gamma)\in\Lambda$, and $K\subset\mathbb{N}$, we define $\delta^{\lambda,\gamma}(K)$ in the following way,
$\delta^{\lambda,\gamma}(K)=\lim_{n\to\infty}\dfrac{|K\cap I_{n}^{\lambda,\gamma}|}{\left(\gamma(n)-\lambda(n)+1\right)},$ (5.2)
where $I_{n}^{\lambda,\gamma}=[\lambda(n),\gamma(n)]$. Because of (5.2), we can state the following lemma (aktuuglu2014korovkin ).
Lemma 6
Let $K$ and $M$ be two subsets of $\mathbb{N}$ and $(\lambda,\gamma)\in\Lambda$. Then the following properties hold true.
  1. (i)
    $\delta^{\lambda,\gamma}(\phi)=0$,
  2. (ii)
    $\delta^{\lambda,\gamma}(\mathbb{N})=0$,
  3. (iii)
    If $K$ is a finite set, then $\delta^{\lambda,\gamma}(K)=0$,
  4. (iv)
    If $K\subset M$, then $\delta^{\lambda,\gamma}(K)\leq\delta^{\lambda,\gamma}(M)$,
  5. (v)
    $\delta^{\lambda,\gamma}(K\cup M)\leq\delta^{\lambda,\gamma}(K)+\delta^{\lambda,\gamma}(M).$
The following definition includes ordinary convergence but along with some well-known matrix methods such as statistical, and $\lambda$-statistical convergence (see aktuuglu2014korovkin ).
Definition 1
A sequence $x$ is said to be $\lambda\gamma$-statistically convergence to $l$ and denoted by $st_{\lambda\gamma}-\lim_{n\to\infty}x_{n}=l$, if for every $\epsilon>0,$
$\delta^{\lambda,\gamma}\left(\{k\in I_{n}^{\lambda,\gamma}:|x_{n}-l|\geq\epsilon\}\right)=\lim_{n\to\infty}\dfrac{\left|\{k\in I_{n}^{\lambda,\gamma}:|x_{n}-l|\geq\epsilon\}\right|}{\left(\gamma(n)-\lambda(n)+1\right)}=0.$
Remark 1
Take $\lambda(n)=1$ and $\gamma(n)=n$. Then $I_{n}^{\lambda,\gamma}=[1,n]$ and
$\delta^{\lambda,\gamma}\left(\{k\in[1,n]:|x_{n}-l|\geq\epsilon\}\right)=\lim_{n\to\infty}\dfrac{\left|\{k\leq n:|x_{n}-l|\geq\epsilon\}\right|}{n}=0.$
Thus for $\lambda(n)=1$ and $\gamma(n)=n$, $\lambda\gamma$-statistical convergence reduces to statistical convergence.
The following two examples show that $\lambda\gamma$-statistical convergence is a non-trivial extension of both ordinary and statistical convergence depending on the choice of $\lambda$ and $\gamma$.
Example 1
Consider the sequence
$c_{n}=\left\{\begin{array}[]{cc}n^{2},&\textrm{ if }n=m^{2}\textrm{ for some }m\in\mathbb{N},\\ \dfrac{1}{n},&\textrm{ otherwise,}\end{array}\right.$ (5.3)
taking $\lambda(n)=1$ and $\gamma(n)=n$, then $\lambda\gamma$-statistical convergence coincide with statistical convergence therefore $st_{\lambda\gamma}-\lim_{n\to\infty}c_{n}=0$.
Example 2
Take $\lambda(n)=2^{2n-1}$, $\gamma(n)=2^{2n}$ and let
$d_{n}=\left\{\begin{array}[]{cc}0,&\textrm{ if }n\in[2^{2m-1},2^{2m}-1]\textrm{ for }m\in\mathbb{N},\\ 1,&\textrm{ otherwise,}\end{array}\right.$ (5.4)
Then the sequence $(d_{n})$ does not converge in statistical sense. Since
|$\dfrac{\left|\{k\in I_{n}^{\lambda,\gamma}:|d_{k}|\geq\epsilon\}\right|}{\left(\gamma(n)-\lambda(n)+1\right)}\leq\dfrac{1}{\left(\gamma(n)-\lambda(n)+1\right)},$
we have $st_{\lambda\gamma}-\lim_{n\to\infty}d_{n}=0$.
Theorem 5
Let $(\lambda,\gamma)\in\Lambda$ and $\mu>0$ and $\beta>0$ be fixed, then
$st_{\lambda\gamma}-\lim_{n}\|W_{n}^{(\beta)}\left(f,\cdot\right)-f\|_{C[0,B]}=0$
holds for every $f\in E$.
Proof: By Lemma 2, we have $\|W_{n}^{(\beta)}\left(1,\cdot\right)-1\|_{C[0,B]}=0$ and $st_{\lambda\gamma}-\|W_{n}^{(\beta)}\left(t,\cdot\right)-t\|_{C[0,B]}=0$.
Again by Lemma 2
$\left\|W_{n}^{(\beta)}(t^{2},\cdot)-t^{2}\right\|_{C[0,B]}\leq\dfrac{x}{n\beta}.$
Now, let $\beta>0$ be a fixed number. For a given $s>0$ choose $\epsilon>0$ such that $\epsilon<s$ and define the following sets:
$V$ ::=:= $\{k\in I_{n}^{\lambda,\gamma}:\|W_{n}^{(\beta)}\left(f,\cdot\right)-f(\cdot)\|_{C[0,B]}\geq s\},$
$V_{1}$ ::=:= $\left\{k\in I_{n}^{\lambda,\gamma}:\dfrac{1}{k}\geq\dfrac{s\beta}{2B}\right\}.$
It is obvious that $V\subset V_{1}$. Therefore, Lemma 6 implies that,
$\delta^{\lambda,\gamma}(V)\leq\delta^{\lambda,\gamma}(V_{1}).$
Using $st_{\lambda\gamma}-\dfrac{1}{n}=0$, we have
$\delta^{\lambda\gamma}(V)=0.$
Hence using (aktuuglu2014korovkin, , Theorem 2), the proof is completed. $\blacksquare$
Now, we consider the following function spaces:
$B_{\rho}:=\{f:\mathbb{R}\to A\subset\mathbb{R}:|f(x)|\leq M_{f}\rho(x)\}$
$C_{\rho}:=\{f:\mathbb{R}\to A\subset\mathbb{R}:f\textrm{ is continuous over }\mathbb{R}\},$
where $\rho(x)$ is a weight function (i.e. $\rho(x)$ is continuous on $\mathbb{R}$ with $\lim_{|x|\to\infty}\rho(x)=\infty$ and $\rho(x)\geq 1$ for all $x\in\mathbb{R}$) and $M_{f}$ is a positive constant depending on $f$. Note that these spaces are Banach spaces endowed with the norm
$\|f\|_{\rho}=\sup_{x\in\mathbb{R}}\dfrac{|f(x)|}{\rho(x)}.$
We will use the following theorem.
Theorem 6 (Duman and Orhan duman2004statistical )
Let $(\lambda,\gamma)\in\Lambda$ and let $\rho_{1}$ and $\rho_{2}$ be weight functions satisfying $\lim_{|x|\to\infty}\dfrac{\rho_{1}(x)}{\rho_{2}(x)}=0$. Assume that $T_{n}$ is a sequence of linear positive operators acting from $C_{\rho_{1}}$ to $B_{\rho_{2}}$. Then for all $f\in C_{\rho_{1}}$,
$st_{\lambda\gamma}-\lim_{n}\|T_{n}(f,\cdot)-f\|_{\rho_{2}}=0$
if and only if
$st_{\lambda\gamma}-\lim_{n}\|T_{n}(F_{v},\cdot)-F_{v}\|_{\rho_{1}}=0,$
where $F_{v}(x)=\dfrac{x^{v}\rho_{1}(x)}{1+x^{2}}(v=0,1,2)$
Choosing the weight functions $\rho_{1}(x)=1+x^{2},\rho_{2}(x)=1+x^{2+\eta}$ for $\eta>0$ and $x\in[0,\infty)$.
Lemma 7
Let $\beta>0$ be fixed, then there exists a constant $N(\beta)>0$ such that
$\dfrac{1}{\rho_{2}(x)}W_{n}^{(\beta)}\left(\rho_{1}(x),x\right)\leq N(\beta)$
holds for all $x\in[0,\infty)$ and $n\in\mathbb{N}$. Furthermore, for all $f\in C_{\rho_{1}}$, we have
$\|W_{n}^{(\beta)}\left(f,\cdot\right)\|\leq N(\beta)\|f\|_{\rho_{1}}.$
Proof: By Lemma 2, we have
$\dfrac{1}{\rho_{2}(x)}W_{n}^{(\beta)}\left(\rho_{1}(t),x\right)$ $=$ $\dfrac{1}{1+x^{2+\mu}}\left[W_{n}^{(\beta)}\left(1,x\right)+W_{n}^{(\beta)}\left(t^{2},\cdot\right)\right]$
$\leq$ $\dfrac{1}{1+x^{2+\mu}}\left[1+x^{2}+\frac{1}{n\beta}x\right]$
$\leq$ $N(\beta).$
To prove the second inequality, we write
$\dfrac{1}{\rho_{2}(x)}\left|W_{n}^{(\beta)}\left(f,x\right)\right|=\dfrac{1}{\rho_{2}(x)}\left|W_{n}^{(\beta)}\left(\rho_{1}\dfrac{f}{\rho_{1}},x\right)\right|\leq\dfrac{\|f\|_{\rho_{1}}}{\rho_{2}(x)}W_{n}^{(\beta)}\left(\rho_{1}(x),x\right)\leq N(\beta)\|f\|_{\rho_{1}}.$ (5.5)
Taking supremum over $x\in[0,\infty)$ in the above inequality, immediately leads to the result. $\blacksquare$
The second inequality in the above lemma shows that $S^{(\beta)}_{n}:C_{\rho_{1}}\to B_{\rho_{2}}$.
Now, we will discussed following weighted $\lambda\beta$-statistical approximation theorem for the Wright operators (1.2):
Theorem 7
Let $(\lambda,\gamma)\in\Lambda$ and $\beta>0$ be fixed, then for each $f\in C_{1+x^{2}}[0,\infty)$ and $\eta>0$, we have
$st_{\lambda\gamma}-\lim_{n}\|W^{(\beta)}_{n}\left(f,\cdot\right)-f\|_{1+x^{2+\gamma}}=0.$
Proof: From Lemma 2, we have
$\|W^{(\beta)}_{n}\left(1,\cdot\right)-1\|_{1+x^{2}}=0.$
Using Lemma 2, we get
$st_{\lambda\gamma}-\|W^{(\beta)}_{n}\left(t,\cdot\right)-x\|_{1+x^{2}}=0.$
Again using Lemma 2, we have
$st_{\lambda\gamma}-\|W^{(\beta)}_{n}\left(t^{2},\cdot\right)-x^{2}\|_{1+x^{2}}$ $=$ $\sup_{x\in[0,\infty)}\dfrac{1}{1+x^{2}}\left[\dfrac{x}{n\beta}\right].$
For a given $r>0$, choose $\epsilon>0$ such that $\epsilon<r$. For fixed $\beta>0$, define sets:
$W$ ::=:= $\{k\in I_{n}^{\lambda,\gamma}:\|W_{n}^{(\beta)}\left(t^{2},\cdot\right)-x^{2}\|_{1+x^{2}}\geq r\}$
$W_{1}$ ::=:= $\left\{k\in I_{n}^{\lambda,\gamma}:\dfrac{1}{k}\geq\dfrac{r\beta}{2}\right\}.$
Since $W\subset W_{1}$, then
$\delta^{\lambda,\gamma}(W)\leq\delta^{\lambda,\gamma}(W_{1}).$
Using $st_{\lambda\gamma}-\dfrac{1}{n}=0$, we have
$\delta^{\lambda,\gamma}(W)=0.$
Finally, applying Theorem 6, hence the proof. $\blacksquare$
Remark 2
As per mention in Remark 1, if $\lambda(n)=1$ and $\gamma(n)=n$, $\lambda\gamma$-statistical convergence reduces to statistical convergence. Therefore, all results obtained here are also valid for statistical convergence.
Remark 3
Let $x_{n}$ be a non-decreasing sequence of positive numbers tending to $\infty$ such that
$x_{n+1}\leq x_{n}+1,\textrm{ and }x_{1}=1.$
Choosing $\lambda(n)=n-x_{n}+1$ and $\gamma(n)=n$, then clearly $I_{n}^{\lambda,\gamma}=[n-x_{n}+1,n]$. Moreover,
$\delta^{\lambda\gamma}\left(\{k:|y_{n}-L|\geq\epsilon\}\right)=\lim_{n\to\infty}\dfrac{\left|\{k\in[n-x_{n}+1,n]:|y_{k}-L|\geq\epsilon\}\right|}{x_{n}}$
this shows that for $\lambda(n)=n-x_{n}+1$ and $\gamma(n)=n$, $\lambda\gamma$-statistical convergence. Therefore, if we take $\lambda(n)=n-x_{n}+1$ and $\gamma_{n}(n)=n$ then all results obtained here are also valid in $x$-statistical sense.
6 Extensions
To obtain an approximation process in the space of integrable functions, we introduce the following new integral modification of Wright operators. Operators in the integral form are more suitable for approximating $L_{p}$ functions with $1\leq p<\infty$. The results for approximating $L_{p}$ functions may have potential applications in machine learning generalization analysis. For the real-valued bounded function $f$ defined on $[0,\infty)$ provided integral exists, we modified the operators (1.2) as follows:
Wright Operators involving Gamma Function:
The generalization of Wright operators involving gamma function is defined as
[…]
Wright Operators involving generalized Gamma Function:
The generalization of Wright operators involving generalized gamma function is defined as
[…]
The approximation properties of operators (6.1) and (6.2) are discussed in patel2024Wright ; patel2024Wright1 .
Kantorovich-type operators:
The Kantorovich-type operators is defined as follows
[…]
Durrmeyer type Modification:
The Durrmeyer type modification of Bernstein operators, Baskakov operators and Szasz–Mirakyan operators are studied in gupta2014convergence ; gupta2021higher ; gupta2021modifications ; ispir2008statistical . Motivated by such type of operators, we generalize the operators $W_{n}^{(\beta)}$ as follows:
[…]
where $b_{n,k}^{[d]}(t)=(-1)^{k}\dfrac{t^{k}}{k!}\xi_{n,d}^{(k)}(t)$. Here $\left\{\xi_{n,d}(x)\right\}$ is a sequence of functions, defined on $[0,b]$, $(b>0)$, and satisfying the following properties for every $n\in\mathbb{N}$ and $k\in\mathbb{N}\cup\{0\}$:
  • 1.
    $\xi_{n,d}\in C^{\infty}\left([a,b]\right),(0\leq a<b);$
  • 2.
    $\xi_{n,d}(0)=1$;
  • 3.
    $\xi_{n,d}$ is completely monotone, that is $(-1)^{k}\xi_{n,d}^{(k)}(x)\geq 0$, $(0\leq x\leq b)$
  • 4.
    there exists an integer $d$ such that $\xi_{n,d}^{(k+1)}(x)=-n\xi_{n+d,d}^{(k)}(x)$, $(n>\max\{0,-d\},x\in[0,b]).$
About these integral operators, we raise the problem to investigate their convergence in $L_{p}$-spaces.
Ethics declarations
Conflict of interest
The authors declares that they have no conflict of interest.
Ethical approval
This article does not contain any studies with human participants or animals performed by the authors.
Informed consent
For this type of study informed consent was not required.
Data availability
Data sharing is not applicable to this article as no new data were created or analysed in this study.