### v2

Abstract
This note aims to present a new sequence of positive linear operators involving the Wright function. Furthermore, the present research established the moments of these newly defined operators and estimated the convergence rate using the classical modulus of continuity. Additionally, the convergence rate in the Lipschitz spaces, their $A$-statistical convergence property have been covered.
keywords:
Wright function; hypergeometric function; Bessel function; $A$-statistical convergence
pacs:
MSC Classification] primary 41A25, 41A30, 41A36
1 Introduction
A new generalization of Bernstein operators was introduced periodically in approximation theory (see [1, 2, 3, 4, 5, 6]). Many positive linear operators are derived from special functions. For instance, the operators were introduced and studied in [7] using the Mittag-Leffler function, the operators were studied in [8] with the help of gamma function, and using Laguerre polynomials, the positive linear operators introduced by Gupta [9]. This note aims to merge two mathematical topics-approximation theory and special functions.
The special function
$\phi(\rho,\beta;z)=\phi_{\rho,\beta}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{k!\Gamma(\rho k+\beta)},~~~(\rho>-1,\beta,z\in\mathbb{C})$ (1.1)
named after the British mathematician E. M. Wright, has appeared for the first time in the case $\rho>0$ in connection with his investigations in the asymptotic theory of partitions in [10].
Note that $\phi(1,1;z)=I_{0}(2\sqrt{z})$,
where $I_{n}(z)=i^{-n}J_{n}(iz)=\dfrac{(z/2)^{n}}{\Gamma(1+n)}{}_{0}F_{1}\left(-;1+n;\dfrac{z^{2}}{4}\right)$, $n$ not a negative integer. The function $I_{n}(z)$ is called the modified Bessel function of the first kind of index $n$. Here, $J_{n}$ is Bessel function and $_{0}F_{1}$ is hypergeometric function. The modified function $I_{n}$ is related to $J_{n}$ in much the same way that the hyperbolic function is related to the trigonometric function.
Also, note that
$\phi_{1,2}(z)=\dfrac{1}{\sqrt{z}}I_{1}\left(2\sqrt{z}\right);~~~~~\phi_{1,m}(z)=\left(\dfrac{1}{\sqrt{z}}\right)^{m/2}I_{m-1}\left(2\sqrt{z}\right).$
It was demonstrated in [11, 12] that, the modified Bessel function $I_{n}(z)$ of the first kind does not have zeros in the right half-plane. [13] contains an analysis of $n$-zeros of the function $I_{n}(z)$ in the left half-plane. Since $I_{n}(z)$ has no zeros in $0,\infty)$, the Wright function $\phi_{1,m}(z)$ also has no zeros in $0,\infty)$.
Let $\beta>1$ be fixed. We provide a novel class of positive linear operators involving the Wright function for every $n\in\mathbb{N}$ as
$W_{n}^{(\beta)}(f;x)=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}f\left(\dfrac{k+\beta}{n}\right)\dfrac{(nx)^{k}}{k!\Gamma(k+\beta)},$ (1.2)
where $f\in E:=\left\{f\in C\left[0,\infty\right):\lim_{x\to\infty}\dfrac{f(x)}{1+x^{2}}\textrm{ is finite }\right\}$ and $C\left[0,\infty\right)$ denote the space of continuous functions defined on $0,\infty)$. Recall that the Banach lattice $E$ is endowed with the norm
$\|f\|_{2}:=\sup_{x\in[0,\infty)}\dfrac{|f(x)|}{1+x^{2}}.$
One can note that the operators $W_{n}^{(\beta)}$ defined in (1.2) are linear and positive. We called it the Wright operators.
Research on approximating continuous signals using a sequence of positive linear operators is ongoing. We provide other findings along the same lines in this note, but with the aid of recently established positive linear operators $W_{n}^{(\beta)}$. There is no prior literature that introduced these operators (1.2). We estimate the moments and central moments of the operators (1.2) up to the fourth order in the second section. The rate of convergence of these positive linear operators is discussed in the next section. In the same section 3, we have proved that $W_{n}^{(\beta)}$ maps $E$ into itself. We have also determined the rate of convergence for these operators by utilizing the modulus of continuity. For the operators $W_{n}^{(\beta)}$, we have established a statistical Voronovskaya-type theorem in section 4.
2 Some Lemmas
To analyze approximation features of the operators (1.2), a few inequalities for the Wright function are required.
Lemma 1.
[14, Theorem 6.1] Let $\alpha,\beta>0$. Then the following assertions are true:
$\Gamma(\beta+\alpha)\phi_{\alpha,\beta+\alpha}(z)\leq\Gamma(\beta)\phi_{\alpha,\beta}(z),\textrm{ for any }z>0.$ (2.1)
Direct calculations allow one to declare the following lemma:
Lemma 2.
Let $\phi_{x}^{2}\left(t\right)=\left(t-x\right)^{2}$, for each $x\geq 0$, $\beta>1$ and $n\in\mathbb{N}$, we have
  1. 1.
    $W_{n}^{(\beta)}(1;x)=1;$
  2. 2.
    $\left|W_{n}^{(\beta)}(t;x)-x\right|\leq\dfrac{\beta}{n};$
  3. 3.
    $\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|\leq\dfrac{x(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}.;$
  4. 4.
    $\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|\leq\dfrac{3x^{2}}{n\beta}+\dfrac{x(1+3\beta+\beta^{2})}{\beta n^{2}}+\dfrac{\beta^{3}}{n^{3}};$
  5. 5.
    $\left|W_{n}^{(\beta)}(t^{4};x)-x^{4}\right|\leq\frac{(4\beta+6)x^{3}}{n\beta(\beta+1)(\beta+2)}+\frac{(6\beta^{2}+12\beta+7)x^{2}}{n^{2}\beta(\beta+1)}+\frac{(4\beta^{3}+6\beta^{2}+4\beta+1)x}{n^{3}\beta}+\frac{\beta^{4}}{n^{4}}.$
Proof.
Since,
$\sum_{k=0}^{\infty}\frac{(nx)^{k}}{k!\Gamma(k+\beta)}=\phi_{1,\beta}(nx),$
one can have $W_{n}^{(\beta)}(1,x)=1$. Using Lemma 1, $\Gamma(\beta+1)\phi_{1,\beta+1}\left(nx\right)\leq\Gamma(\beta)\phi_{1,\beta}\left(nx\right)$ for any $x\in[0,\infty)$ and $n\in\mathbb{N}$, we get
$W_{n}^{(\beta)}(t;x)$ $=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{1}{n}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}+\dfrac{\beta}{n}$
$=\dfrac{x}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{(nx)^{k}}{k!\Gamma(k+\beta+1)}+\dfrac{\beta}{n}=x\dfrac{\phi_{1,\beta+1}(nx)}{\phi_{1,\beta}(nx)}+\dfrac{\beta}{n}\leq x+\dfrac{\beta}{n}\textrm{ for }\beta>1.$
Again applying Lemma 1, $\Gamma(\beta+2)\phi_{1,\beta+2}\left(nx\right)\leq\Gamma(\beta+1)\phi_{1,\beta+1}\left(nx\right)\leq\Gamma(\beta)\phi_{1,\beta}\left(nx\right)$ for any $x\in[0,\infty)$ and $n\in\mathbb{N}$, we get
$W_{n}^{(\beta)}(t^{2};x)$ $=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=1}^{\infty}\dfrac{(k+\beta)^{2}}{n^{2}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}$
$=\frac{1}{n^{2}\,\phi_{1,\beta}(nx)}\Bigg[(nx)^{2}\sum_{k=0}^{\infty}\frac{(nx)^{k}}{k!\,\Gamma(k+2+\beta)}+(1+2\beta)\,nx\sum_{m=0}^{\infty}\frac{(nx)^{m}}{m!\,\Gamma(m+1+\beta)}+\beta^{2}\,\phi_{1,\beta}(nx)\Bigg]$
$=\frac{1}{n^{2}\,\phi_{1,\beta}(nx)}\Big[(nx)^{2}\,\phi_{1,\beta+2}(nx)+(1+2\beta)\,nx\,\phi_{1,\beta+1}(nx)+\beta^{2}\,\phi_{1,\beta}(nx)\Big].$
Since $\beta>1$,
$\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|\leq\dfrac{x(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}.$
Now,
$W_{n}^{(\beta)}(t^{3};x)$ $=$ $\frac{1}{n^{3}\phi_{1,\beta}(nx)}\sum_{k=0}^{\infty}(k+\beta)^{3}\frac{(nx)^{k}}{k!,\Gamma(k+\beta)}$
$=$ $\frac{1}{n^{3}\phi_{1,\beta}(nx)}\Big[(nx)^{3}\phi_{1,\beta+3}(nx)+3(1+\beta)(nx)^{2}\phi_{1,\beta+2}(nx)$
$+(1+3\beta+3\beta^{2})nx\phi_{1,\beta+1}(nx)+\beta^{3}\phi_{1,\beta}(nx)\Big].$
$=$ $x^{3}\dfrac{\phi_{1,\beta+3}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}+3(1+\beta)x^{2}\dfrac{\phi_{1,\beta+2}\left(nx\right)}{n\phi_{1,\beta}\left(nx\right)}+x(1+3\beta+\beta^{2})\dfrac{\phi_{1,\beta+1}\left(nx\right)}{n^{2}\phi_{1,\beta}\left(nx\right)}+\dfrac{\beta^{3}}{n^{3}}$
$\leq$ $\left|\dfrac{x^{3}\Gamma(\beta)}{\Gamma(\beta+3)}\right|+\left|\dfrac{3(1+\beta)x^{2}\Gamma(\beta)}{n\Gamma(\beta+2)}\right|+\left|\dfrac{x(1+3\beta+\beta^{2})\Gamma(\beta)}{n^{2}\Gamma(\beta+1)}\right|$
$=$ $\dfrac{x^{3}}{\beta(\beta+1)(\beta+2)}+\dfrac{3x^{2}}{n\beta}+\dfrac{x(1+3\beta+\beta^{2})}{\beta n^{2}}+\dfrac{\beta^{3}}{n^{3}}.$
Since, $\beta>1$,
$\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|\leq\dfrac{3x^{2}}{n\beta}+\dfrac{x(1+3\beta+\beta^{2})}{\beta n^{2}}+\dfrac{\beta^{3}}{n^{3}}.$
Similarly,
$W_{n}^{(\beta)}(t^{4};x)$ $=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{(k+\beta)^{4}}{n^{4}}\dfrac{(nx)^{k}}{k!\Gamma(k+\beta)}$
$=$ $\frac{1}{n^{4}\phi_{1,\beta}(nx)}\Big[(nx)^{4}\phi_{1,\beta+4}(nx)+(4\beta+6)(nx)^{3}\phi_{1,\beta+3}(nx)$
$+(6\beta^{2}+12\beta+7)(nx)^{2}\phi_{1,\beta+2}(nx)+(4\beta^{3}+6\beta^{2}+4\beta+1)nx\phi_{1,\beta+1}(nx)+\beta^{4}\phi_{1,\beta}(nx)\Big].$
$=$ $x^{4}\dfrac{\phi_{1,\beta+4}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}+(4\beta+6)x^{3}\dfrac{\phi_{1,\beta+3}\left(nx\right)}{n\phi_{1,\beta}\left(nx\right)}+(6\beta^{2}+12\beta+7)x^{2}\dfrac{\phi_{1,\beta+2}\left(nx\right)}{n^{2}\phi_{1,\beta}\left(nx\right)}$
$+(4\beta^{3}+6\beta^{2}+4\beta+1)x\dfrac{\phi_{1,\beta+1}\left(nx\right)}{n^{3}\phi_{1,\beta}\left(nx\right)}+\dfrac{\beta^{4}}{n^{4}}$
$\leq$ $\left|\dfrac{x^{4}\Gamma(\beta)}{\Gamma(\beta+4)}\right|+\left|\dfrac{(4\beta+6)x^{3}\Gamma(\beta)}{n\Gamma(\beta+3)}\right|+\left|\dfrac{(6\beta^{2}+12\beta+7)x^{2}\Gamma(\beta)}{n^{2}\Gamma(\beta+2)}\right|$
$+\left|\dfrac{(4\beta^{3}+6\beta^{2}+4\beta+1)x\Gamma(\beta)}{n^{3}\Gamma(\beta+1)}\right|+\dfrac{\beta^{4}}{n^{4}}$
$=$ $\dfrac{x^{4}}{\beta(\beta+1)(\beta+2)(\beta+3)}+\dfrac{(4\beta+6)x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{(6\beta^{2}+12\beta+7)x^{2}}{n^{2}\beta(\beta+1)}$
$+\dfrac{(4\beta^{3}+6\beta^{2}+4\beta+1)x}{n^{3}\beta}+\dfrac{\beta^{4}}{n^{4}}.$
Since, $\beta>1$,
$\left|W_{n}^{(\beta)}(t^{4};x)-x^{4}\right|\leq\dfrac{(4\beta+6)x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{(6\beta^{2}+12\beta+7)x^{2}}{n^{2}\beta(\beta+1)}+\dfrac{(4\beta^{3}+6\beta^{2}+4\beta+1)x}{n^{3}\beta}+\dfrac{\beta^{4}}{n^{4}}.$
Lemma 3.
Let $\phi_{x}^{i}\left(t\right)=\left(t-x\right)^{i}$, $i\in\mathbb{N}$, for each $x\geq 0$, $\beta>1$ and $n\in\mathbb{N}$, we have
  1. 1.
    $W_{n}^{(\beta)}(\phi_{x}^{1};x)\leq\dfrac{\beta}{n}$
  2. 2.
    $W_{n}^{(\beta)}(\phi_{x}^{2};x)\leq\frac{x(1+2\beta)}{\beta n}+\frac{2x\beta}{n}+\frac{\beta^{2}}{n^{2}}.$
  3. 3.
    $W_{n}^{(\beta)}(\phi_{x}^{3};x)\leq\frac{3x^{2}}{n\beta}+\frac{x(1+3\beta+\beta^{2})}{\beta n^{2}}+\frac{\beta^{3}}{n^{3}}+3x\left(\frac{x(1+2\beta)}{\beta n}+\frac{\beta^{2}}{n^{2}}\right)+\frac{3x^{2}\beta}{n}.$
  4. 4.
    $W_{n}^{(\beta)}(\phi_{x}^{4};x)\leq\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta\right)\frac{x^{3}}{n}+\left(\frac{6\beta^{2}+12\beta+7}{\beta(\beta+1)}+\frac{4}{\beta}+12+4\beta+6\beta^{2}\right)\frac{x^{2}}{n^{2}}+\left(4\beta^{3}+4\beta^{2}+6\beta+4+\frac{1}{\beta}\right)\frac{x}{n^{3}}+\frac{\beta^{4}}{n^{4}}.$
Proof.
Using Lemma 2 and simple computation gives $W_{n}^{(\beta)}(\phi_{x}^{1};x)=0$ and
$W_{n}^{(\beta)}(\phi_{x}^{2};x)\leq\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|+2x\left|W_{n}^{(\beta)}(t;x)-x\right|+x^{2}\left|W_{n}^{(\beta)}(1;x)-1\right|\leq\frac{x(1+2\beta)}{\beta n}+\frac{2x\beta}{n}+\frac{\beta^{2}}{n^{2}}.$
Further,
$W_{n}^{(\beta)}(\phi_{x}^{3};x)$ $\leq\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|+3x\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|+3x^{2}\left|W_{n}^{(\beta)}(t;x)-x\right|+x^{3}\left|W_{n}^{(\beta)}(1;x)-1\right|$
$\leq\dfrac{3x^{2}}{n\beta(\beta+1)}+\dfrac{x}{n^{2}\beta}+3x\left(\dfrac{x}{n\beta}\right)=\dfrac{3x^{2}(\beta+2)}{n\beta(\beta+1)}+\dfrac{x}{n^{2}\beta}.$
Finally, we have
$W_{n}^{(\beta)}(\phi_{x}^{4};x)$ $\leq\left|W_{n}^{(\beta)}(t^{4};x)-x^{4}\right|+4x\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|$
$+6x^{2}\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|+4x^{3}\left|W_{n}^{(\beta)}(t;x)-x\right|+x^{4}\left|W_{n}^{(\beta)}(1;x)-1\right|$
$\leq\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{12}{\beta}+\frac{6(1+2\beta)}{\beta}+4\beta\right)\frac{x^{3}}{n}$
$+\left(\frac{6\beta^{2}+12\beta+7}{\beta(\beta+1)}+\frac{4(1+3\beta+\beta^{2})}{\beta}+6\beta^{2}\right)\frac{x^{2}}{n^{2}}+\left(\frac{4\beta^{3}+6\beta^{2}+4\beta+1}{\beta}+4\beta^{3}\right)\frac{x}{n^{3}}+\frac{\beta^{4}}{n^{4}}.$
3 Rate of Convergence
In this section, the rate of convergence of $W_{n}^{(\beta)}$ is discussed. The following lemma proves that $W_{n}^{(\beta)}$ maps $E$ into itself.
Lemma 4.
Let $\beta>1$ be fixed, then there exists a constant $M(\beta)$ such that,
$\omega(x)W_{n}^{(\beta)}\left(\dfrac{1}{\omega},x\right)\leq M(\beta)$
holds for all $x\in[0,\infty)$, $n\in\mathbb{N}$ and $\omega(x)=\dfrac{1}{1+x^{2}}$. Furthermore, for all $f\in E$, we have
$\|D_{n}^{(\beta)}(f,\cdot)\|_{2}\leq M(\beta)\|f\|_{2}$
Proof.
From Lemma 2, we have
$\omega(x)W_{n}^{(\beta)}\left(\dfrac{1}{\omega},x\right)$ $=$ $\dfrac{1}{1+x^{2}}\left[W_{n}^{(\beta)}\left(1,x\right)+W_{n}^{(\beta)}\left(t^{2},x\right)\right]$
$\leq$ $\dfrac{1}{1+x^{2}}\left[1+x^{2}+\dfrac{x(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}\right]$
$\leq$ $M(\beta)$
This follows, following inequality
$\omega(x)\left|W_{n}^{(\beta)}\left(f,x\right)\right|=\omega(x)\left|W_{n}^{(\beta)}\left(\omega\dfrac{f}{\omega},x\right)\right|\leq\|f\|_{2}\omega(x)W_{n}^{(\beta)}\left(\dfrac{1}{\omega},x\right)\leq M(\beta)\|f\|_{2}$
Taking supremum over $x\in[0,\infty)$ in the above inequality, gives the result. ∎
Now, recall that the usual modulus of continuity of $f$ on the closed interval $[0,B]$ is defined by
$\omega_{B}(f,\delta)=\sup\{|f(t)-f(x)|:|t-x|\leq\delta,~~x,t\in[0,B]\}$
It is well known that, for a function $f\in E$, we have $\lim_{\delta\to\infty}\omega_{B}(f,\delta)=0$. The next theorem gives the rate of convergence of the operators $W_{n}^{(\beta)}(f,x)$, for all $f\in E$.
Theorem 1.
Let $\beta>1$, $f\in E$ and $\omega_{B+1}(f,\delta)$, $(B>0)$ be its modulus of continuity on the finite interval $0,B+1]\subset[0,\infty)$, then
$\|W_{n}^{(\beta)}\left(f,\cdot\right)-f(\cdot)\|_{C[0,B]}\leq M_{f}(\beta,B)\delta_{n}(\beta,B)+2\omega_{B+1}\left(f,\delta_{n}^{1/2}(\beta,B)\right)$
where $\delta_{n}(\beta,B)=\dfrac{B(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}$ and $M_{f}(\beta,B)$ is an absolute constant depending on $f$, $\beta$ and $B$.
Proof.
Let $\beta>0$ be fixed. For $x\in[0,B]$ and $t\leq B+1$, we have following well-known inequality
$|f(t)-f(x)|\leq\omega_{B+1}(f,|t-x|)\leq\left(1+\dfrac{|t-x|}{\delta}\right)\omega_{B+1}(f,\delta)$ (3.1)
where $\delta>0$. Now, for $x\in[0,B]$ and $t>B+1$, using the fact that tx>1t-x>1, we have
$|f(t)-f(x)|$ $\leq$ $A_{f}(1+x^{2}+t^{2})$ (3.2)
$\leq$ $A_{f}(2+3x^{2}+2(t-x)^{2})$
$\leq$ $6A_{f}(1+B^{2})(t-x)^{2}$
Using (3.1) and (3.2), we get for all $x\in[0,B]$ and $t\geq 0$, we get
$|f(t)-f(x)|$ $\leq$ $6A_{f}(1+B^{2})(t-x)^{2}+\left(1+\dfrac{|t-x|}{\delta}\right)\omega_{B+1}(f,\delta)$ (3.3)
Therefore,
$\left|W_{n}^{(\beta)}\left(f,x\right)-f(x)\right|\leq 6A_{f}(1+B^{2})W_{n}^{(\beta)}\left(\phi_{x}^{2},x\right)+\left(1+\dfrac{W_{n}^{(\beta)}\left(|t-x|,x\right)}{\delta}\right)\omega_{B+1}(f,\delta)$
Applying Cauchy-Schwarz inequality and Lemma 2, we get
$\left|W_{n}^{(\beta)}\left(f,x\right)-f(x)\right|$ $\leq$ $6A_{f}(1+B^{2})W_{n}^{(\beta)}\left(\phi_{x}^{2},x\right)+\left(1+\dfrac{\left[W_{n}^{(\beta)}\left(\phi_{x}^{2},x\right)\right]^{1/2}}{\delta}\right)\omega_{B+1}(f,\delta)$
$\leq$ $6A_{f}(1+B^{2})\left(\dfrac{B(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}\right)+\left(1+\dfrac{\left[\dfrac{B(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}\right]^{1/2}}{\delta}\right)\omega_{B+1}(f,\delta)$
$\leq$ $M_{f}(\beta,B)\delta_{n}^{2}(\beta,B)+2\omega_{B+1}\left(f,\left(\delta_{n}(\beta,B)\right)^{1/2}\right)$
where $M_{f}(\beta,B)=6A_{f}(1+B^{2})$ and $\delta_{n}(\beta,B)=\dfrac{B(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}$. Hence the proof. ∎
4 $A$-statistical Convergence
In this section, first we discussed some definitions and notations on the concept of $A$-statistical convergence. Let $A=\left(a_{nk}\right)$, $\left(n,k\in\mathbb{N}\right)$, be a non-negative, infinite summability matrix. For a given sequence $x:=(x_{k})$, the $A$-transform of $x$ denoted by $Ax:\left((Ax)_{n}\right)$ is defined as
$\left(Ax\right)_{n}=\sum_{k=1}^{\infty}a_{nk}x_{k},$
provided the series converges for each $n$. AA is said to be regular if $\lim_{n}(Ax)_{n}=L$ whenever $\lim_{n}x_{n}=L$. We say that, the sequence $x=(x_{n})$ is $A$-statistically convergent to $L$ and write $st_{A}-\lim_{n}x_{n}=L$ if for every $\epsilon>0$, $\lim_{n}\sum_{k:|x_{k}-L|\geq\epsilon}a_{nk}=0$.
Replacing $A$ by $C_{1}$, the Cesáro matrix of order one, the $A$-statistical convergence reduces to statistical convergence. Similarly, if we take $A=I$, the identity matrix, then $A$-statistical convergence coincides with ordinary convergence. The statistical convergence of various types of operators has been studied by several researchers (see [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]). Now, we prove weighted Korovkin theorem via $A$-statistical convergence.
Here, we recall the weighted Korovkin type approximation theorem for the A-statistical convergence was given by Duman and Orhan in [30].
Theorem 2.
[30] Let $A$ be a non-negative regular summability matrix and let $\bar{\rho}_{1}$; $\bar{\rho}_{2}$ weight functions such that
$\lim_{|x|\to\infty}\dfrac{\bar{\rho}_{1}(x)}{\bar{\rho}_{2}(x)}=0.$
Assume that $(T_{n})_{n\geq 1}$ is a sequence of positive linear operators from $C_{\bar{\rho}_{1}}(\mathbb{R})$ into $B_{\bar{\rho}_{2}}(\mathbb{R})$, One has
$st_{A}-\lim_{n}\|T_{n}f-f\|_{\bar{\rho}_{2}}=0,$
for all $f\in C_{\bar{\rho}_{1}}(\mathbb{R})$ if and only if
$st_{A}-\lim_{n}\|T_{n}F_{v}-F_{v}\|_{\bar{\rho}_{2}}=0,\textrm{ for all }v=0,1,2,$
where
$F_{v}(x)=\dfrac{x^{v}\bar{\rho}_{1}(x)}{1+x^{2}},v=0,1,2.$
By using this theorem the following Korovkin type theorem can be proved for $\left(W_{n}^{(\beta)}\right)$.
Theorem 3.
Let $A=(a_{nk})$ be a non-negative regular summability matrix, $\beta>1$ be fixed and $x\in[0,\infty)$, then for all $f\in E$, we have
$st_{A}-\lim_{n}\|W_{n}^{(\beta)}\left(f,\cdot\right)-f\|_{2}=0$
Proof.
From [30, p. 191, Th. 3], it is sufficient to show that $st_{A}-\lim_{n}\|W_{n}^{(\beta)}\left(t^{i},\cdot\right)-x^{i}\|_{2}=0$, where $i=0,1,2$.
In view of Lemma 2, it follows that
$st_{A}-\lim_{n}\|W_{n}^{(\beta)}\left(1,\cdot\right)-1\|_{2}=0$
and
$st_{A}-\lim_{n}\|W_{n}^{(\beta)}\left(t,\cdot\right)-x\|_{2}=0.$
Now,
$\|W_{n}^{(\beta)}\left(t^{2},\cdot\right)-x^{2}\|_{2}\leq\sup_{x\geq 0}\left(\dfrac{x^{2}(1+2\beta)}{\beta n(1+x^{2})}+\dfrac{x\beta^{2}}{n^{2}(1+x^{2})}\right)\leq\dfrac{1+2\beta}{n\beta}+\dfrac{\beta^{2}}{n^{2}}$
Given $r>0$, choose $\epsilon>0$ such that $\epsilon<r$. For fixed $\beta>1$, define the following sets:
$U:=\left\{n:\|W_{n}^{(\beta)}\left(t,\cdot\right)-x\|_{2}\geq\epsilon\right\}$
$U_{1}:=\left\{n:\dfrac{(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}\geq\dfrac{\epsilon}{2}\right\}$
Then it is clear that $U\subset U_{1}$, this gives
$\sum_{k\in U}a_{jk}\leq\sum_{k\in U_{1}}a_{jk}$ (4.1)
Letting $j\to\infty$ in (4.1), we have $\lim_{j}\sum_{k\in U}a_{jk}=0$. This proves that $st_{A}-\lim_{n}\dfrac{1}{n\beta}=0$, this also implies
$st_{A}-\lim_{n}\omega_{B+1}\left(f,\sqrt{\dfrac{1}{n\beta}}\right)=0$
Using theorem 1, we get desired result.∎
Lemma 5.
Let $A=(a_{in})$ be a non-negative regular summability matrix, then we have
$st_{A}-\lim_{n\to\infty}nW_{n}^{(\beta)}\left((t-x)^{4},x\right)=\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta.$
uniformly with respect to $x\in[0,b]$ with $b>0$.
Proof.
Note that
$W_{n}^{(\beta)}\left((t-x)^{4},x\right)$ $\leq$ $\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta\right)\frac{x^{3}}{n}+\left(\frac{6\beta^{2}+12\beta+7}{\beta(\beta+1)}+\frac{4}{\beta}+12+4\beta+6\beta^{2}\right)\frac{x^{2}}{n^{2}}$
$+\left(4\beta^{3}+4\beta^{2}+6\beta+4+\frac{1}{\beta}\right)\frac{x}{n^{3}}+\frac{\beta^{4}}{n^{4}}.$
This gives
$\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta\right)\right|\leq\dfrac{h_{1}(\beta)}{n}x^{2}+\dfrac{h_{2}(\beta)}{n^{2}}x+\dfrac{\beta^{4}}{n^{3}}$
where $h_{1}$ and $h_{2}$ are some functions of $\beta$.
For $x\in[0,b]$, we have
$\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta\right)\right|\leq B\left\{\dfrac{1}{n}+\dfrac{1}{n^{2}}+\dfrac{1}{n^{3}}\right\}$ (4.2)
where $B=\max\{h_{1}(\beta)b^{2},h_{2}(\beta)b,\beta^{4}\}$. Now, for a given $\epsilon>0$, define the following sets:
$D$ $:=$ $\left\{n:\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta\right)\right|\geq\epsilon\right\}$
$D_{1}$ $:=$ $\left\{n:\dfrac{1}{n}\geq\dfrac{\epsilon}{2B}\right\}$
$D_{2}$ $:=$ $\left\{n:\dfrac{1}{n}\geq\sqrt{\dfrac{\epsilon}{2B}}\right\}$
$D_{3}$ $:=$ $\left\{n:\dfrac{1}{n}\geq\sqrt[3]{\dfrac{\epsilon}{B}}\right\}$
Hence, by inequality (4.2), we see that $D\subset D_{1}\cup D_{2}\cup D_{3}$. Then for any $j\in\mathbb{N}$, we have
$\sum_{n\in D}a_{jn}\leq\sum_{n\in D_{1}}a_{jn}+\sum_{n\in D_{2}}a_{jn}\sum_{n\in D_{3}}a_{jn}.$ (4.3)
Taking limit as $j\to\infty$ on the both sides of above inequality and using the fact that $st_{A}-\lim_{n\to\infty}\left(\dfrac{1}{n}\right)=0$, we conclude that
$\lim_{j\to\infty}\sum_{n\in D}a_{jn}=0.$
Hence the result. ∎
Theorem 4 (Statistical Voronovskaya-type theorem for the operator $W_{n}^{(\beta)}$).
Let $A=(a_{jn})$ be a nonnegative regular summability matrix, $\beta>1$ then for every $f\in E$ with $f^{\prime},f^{\prime\prime}\in E$, we have
$st_{A}-\lim_{n\to\infty}n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)=\frac{1+2\beta+2\beta^{2}}{\beta}\dfrac{x}{2}f^{\prime\prime}(x)$
uniformly with respect to $x\in[0,b]$ with $b>0$.
Proof.
Let $f,f^{\prime},f^{\prime\prime}\in E$ and $x\in[0,b]$. Define the function $\Phi_{x}$ by
$\Phi_{x}(t)=\left\{\begin{array}[]{ll}\dfrac{f(t)-f(x)-(t-x)f^{\prime}(x)-\frac{1}{2}(t-x)^{2}f^{\prime\prime}(x)}{(t-x)^{2}},&\textrm{ if }t\neq x\\ 0,&\textrm{ if }t=x\end{array}\right$.
Then, it is clear that $\Phi_{x}(x)=0$. Also, observe that the function $\Phi_{x}(\cdot)$ belongs to $E$. Hence, by Taylor’s theorem, we get
$f(t)=f(x)+(t-x)f^{\prime}(x)+\dfrac{(t-x)^{2}}{2}f^{\prime\prime}(x)+(t-x)^{2}\Phi_{x}(t)$
Now the definition of the operators (1.2) implies that
$W_{n}^{(\beta)}(f,x)-f(x)=f^{\prime}(x)W_{n}^{(\beta)}(t-x,x)+\dfrac{1}{2}f^{\prime\prime}(x)W_{n}^{(\beta)}((t-x)^{2},x)+W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)$
Therefore, using lemma 2, we have
$\left|n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)-\frac{1+2\beta+2\beta^{2}}{2\beta}\,xf^{\prime\prime}(x)\right|$ $\leq$ $n\left|W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)\right|$ (4.4)
Applying the Cauchy-Schwarz inequality to the second term on the right-hand side of (4.4), then we see that
$\left|W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)\right|\leq\sqrt{W_{n}^{(\beta)}((t-x)^{4},x)}\sqrt{W_{n}^{(\beta)}(\Phi_{x}^{2}(t),x)}$
this yields
$n\left|W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)\right|\leq\sqrt{n^{2}W_{n}^{(\beta)}((t-x)^{4},x)}\sqrt{W_{n}^{(\beta)}(\Phi_{x}^{2}(t),x)}$ (4.5)
Let $\eta_{x}(t):=\Phi_{x}^{2}(t)$. In this case, observe that $\eta_{x}(x)=0$ and $\eta_{x}(\cdot)\in E$. Then it follows from Theorem 3 that
$st_{A}-\lim_{n\to\infty}W_{n}^{(\beta)}(\Phi_{x}^{2}(t),x)=st_{A}-\lim_{n\to\infty}W_{n}^{(\beta)}(\eta_{x}(t),x)=\eta_{x}(x)=0$ (4.6)
uniformly with respect to $x\in[0,b]$. Now considering (4.5) and (4.6), and also Lemma 5, we immediately lead to
$st_{A}-\lim_{n\to\infty}n\left(W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)\right)=0$ (4.7)
uniformly with respect to $x\in[0,b]$. Using (4.4) to (4.7) and also considering $st_{A}-\lim_{n\to\infty}\dfrac{1}{n}=0$, we have
$st_{A}-\lim_{n\to\infty}n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)\frac{1+2\beta+2\beta^{2}}{2\beta}\,xf^{\prime\prime}(x)$
uniformly with respect to $x\in[0,b]$. ∎
Ethics declarations
Conflict of interest
The authors declares that they have no conflict of interest.
Ethical approval
This article does not contain any studies with human participants or animals performed by the authors.
Informed consent
For this type of study informed consent was not required.
Data availability
Data sharing is not applicable to this article as no new data were created or analysed in this study.